OpenStudy (anonymous):

What is the solution to the equation 14 x + 3 = 21 x ≈ –2.133 x ≈ –1.846 x ≈ 3.867 x ≈ 4.154

OpenStudy (anonymous):

None of the choices.

OpenStudy (anonymous):

Ya I didn't get anything that was given.

OpenStudy (anonymous):

14x=21-3 14x=18 x=1.2857...

OpenStudy (anonymous):

Sorry its supposed to be 14 ^x+3 = 21

OpenStudy (anonymous):

thanks